Find the distance between the lines with the equations y=(2/7)x+4 and y=(2/7)x-2
Answer by Alan3354(69443)

Find the distance between the lines with the equations y=(2/7)x+4 and y=(2/7)x-2
-----------------
The distance between the 2 lines on the y-axis = 6
The angle between the x-axis and the lines is the atan of the slope,
= atan(2/7)
------
Distance between the 2 lines is
6*cos(atan(2/7))
distane =~ 5.769 units